OpenStudy (anonymous):

what is the sum of the measures of the interior angles of a convex 16-gon A.1800 B. 2340 C. 2520 D. 2700 E. 2800 PLEASE HELP :) THANKS

OpenStudy (amistre64):

360/16 maybe?

OpenStudy (amistre64):

180-(360/16)

OpenStudy (amistre64):

getting warmer ..

OpenStudy (amistre64):

16(180-(360/16))

OpenStudy (amistre64):

2520 if i did it right lol

OpenStudy (anonymous):

(1-2/n)*180 is one internal angle where n is the number of sides. sum of n sides (n-2)*180 14*180=2520

OpenStudy (anonymous):

amistre is right

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

yeah the formula is (n-2)*180 (16-2)*180 = 14*180 =2520.

OpenStudy (anonymous):

360/16 = 22.5 180 - 22.5 = 157.5 16 x 157.5 = 2520

OpenStudy (amistre64):

theres a formula? :)

OpenStudy (amistre64):

why am I having to do these on the fly like a white guy then?

OpenStudy (anonymous):

yes, there is a formula.

OpenStudy (anonymous):

YES THERE IS A FORMULA TO THE QUESTION

OpenStudy (amistre64):

im to old to commit resources to stuff im just gonna forget ;)